# Theory of Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Using Super-Resolution Operators as Background Error Covariance Matrices

### v3

@@ III.2 Reduction of the ELBO to the 3D-Var objective function
where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from the inputs $\bm{x}$ and $\bm{y}$. The function $l_{\text{3D-Var}}$ is equivalent to the 3D-Var objective function,Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018); Lorenc (1986) and the vector $\bm{a}$ is obtained by minimizing $l_{\text{3D-Var}}$. The first term of $l_{\text{3D-Var}}$, which comes from the reconstruction error in Eq. (2c), compares the observation vector $\bm{y}$ and the analysis $\bm{a}$ through the observation operator $\mathsf{H}$. The second term of $l_{\text{3D-Var}}$, which comes from the KL divergence in Eq. (2c), measures the difference between the analysis $\bm{a}$ and the background state $\bm{x}$. By minimizing $l_{\text{3D-Var}}$, $\bm{y}$ and $\bm{x}$ are integrated, and the true state is estimated as $\bm{a}$. These results suggest that the ELBO describes the balance between observations and background states in estimating the true state.
The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is linear (i.e., $\mathsf{H}=H$, where $H$ is a matrix).Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018)
@@
where $K$ is known as the Kalman gain matrix. The second term in Eq. (9a) describes the non-local nature of DA. The background state $\bm{x}$ is corrected through the spatially sparse observations $\bm{y}$ by the non-diagonal matrix $K$. Typically, this non-diagonal property originates from the non-zero off-diagonal elements of the background error covariance matrix $B$. In other words, the non-local correction is realized through the spatial correlation patterns of background values. The necessity of spatial correlations leads to a difficulty for DA by neural networks (Section III.3).
III.3 Implementation of DA by CVAE
DA can be performed by solving the minimization problem of $l_{\text{Gauss}}$. However, this minimization at every assimilation time step would be computationally intensive. We propose to efficiently conduct this minimization using a trained CVAE.
Generally, the loss function for VAEs, including CVAEs, is an ELBO.Kingma and Welling (2014); Rezende, Mohamed, and Wierstra (2014); Kingma et al. (2014); Sohn, Lee, and Yan (2015); Kingma and Welling (2019) This can be understood from the ELBO formula, Eq. (7). To evaluate $l_{\text{Gauss}}$, the analysis $\bm{a}$ and its covariance $V$ are first computed from a given input of $\bm{x}$ and $\bm{y}$; then, the observation vector $\bm{\hat{y}}$ ($=\mathsf{H}(\bm{\hat{z}})$) is reconstructed from $\bm{a}$ and $V$ via Eq. (5). The value of $l_{\text{Gauss}}$ decreases as $\bm{\hat{y}}$ becomes closer to the input $\bm{y}$. The structure that reconstructs the input is common to autoencoders.Dong et al. (2018) The second and third terms in Eq. (7) are considered as the conditions for the estimated true state $\bm{\hat{z}}$. For instance, $l_{\text{Gauss}}$ is decreased as $\bm{a}$ becomes close to $\bm{x}$. Thus, a CVAE is necessary to incorporate these conditions, and this CVAE is trained by minimizing $l_{\text{Gauss}}$.
The CVAE consists of an encoder and a decoder (Fig. 2). The encoder estimates $\bm{a}$ and $V$ from the input $\bm{x}$ and $\bm{y}$, while the decoder reconstructs $\bm{\hat{y}}$. In other words, the encoder performs assimilation and the decoder acts as the nonlinear observation operator. The encoder learns how to integrate $\bm{x}$ and $\bm{y}$, yielding the analysis $\bm{a}$, while the decoder learns how to convert $\bm{a}$ into the observation space. Once the CVAE is trained, the encoder performs DA without minimizing $l_{\text{Gauss}}$ as long as the statistics of the background $\bm{x}$ and the observations $\bm{y}$ do not change significantly.
@@ III.3 Implementation of DA by CVAE
The covariance matrices $R$ and $B$ in $l_{\text{Gauss}}$ are treated as prescribed parameters (i.e., hyperparameters). Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices: $R=rI_{n_{y}}$ and $B=bI_{n_{z}}$, where $I_{n}$ is an identity matrix of size $n$, and $r$ and $b$ are positive real parameters. However, off-diagonal elements of $B$ describe the spatial correlation patterns of background states and are essential for the non-local estimation in DA (Section III.2).
There are several methods for implementing non-local estimation. For instance, latent variables can be transformed to have covariance or even follow non-Gaussian distributions by combining VAEs and normalizing flows,Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) where this transformation is learnable from data. An alternative approach is to divide observed values used as input to CVAEs and those targeted for reconstruction. CVAEs reconstruct observations from the other observations at different grid points, which enables CVAEs to learn non-locality. A similar technique is employed in MetNet-3, a surrogate neural network for weather predictions developed by Google.Andrychowicz et al. (2023) As a final method, we adopt super-resolution (SR). The non-locality is incorporated into CVAEs through SR even when $B$ and $R$ are diagonal.
IV SRDA using the ELBO
We incorporate SR into the DA theory proposed in Section III. A key characteristic of SR models is their non-locality:Dong et al. (2014); Gu and Dong (2021) SR models map LR patterns with a certain spatial extent to HR patterns. These SR mappings (i.e., SR operators) are integrated into DA, resulting in SRDA, which facilitates the non-local correction of background values by observations across distant grid points.
IV.1 Integration of SR into DA
To develop a theory of SRDA, we specify the resolutions of a background state $\bm{x}$, observations $\bm{y}$, and the true state $\bm{z}$: $\bm{x}$ is defined at LR grid points, while $\bm{y}$ and $\bm{z}$ are defined at HR grid points. The vector forms are obtained by arranging three-dimensional grid point values into columns. Observations are not available at all locations due to various factors, such as measurement errors. These missing grid points are excluded when arranging observations into the column vector $\bm{y}$.
@@ IV.1 Integration of SR into DA
$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\exp\left\{-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\right\},$ (10a)
@@
where each variance $v_{i}(\bm{x},\bm{y})$ ($i=1,\ldots,n_{z}$) is a positive value determined by the vectors $\bm{x}$ and $\bm{y}$.
There are two essential changes from Eqs. (3) to (10). First, an SR operator $\mathsf{F}$ is introduced in Eq. (10b), where $\mathsf{F}$ maps an $n_{x}$-dimensional vector into an $n_{z}$-dimensional vector. Since $\bm{x}$ and $\bm{z}$ are defined at LR and HR grid points, respectively (i.e., $n_{x}<n_{z}$), this mapping performs SR. Comparing Eqs. (3b) and (10b), the mean of the prior distribution $p\left(\bm{z}\mid\bm{x}\right)$ is replaced with the super-resolved background state $\mathsf{F}(\bm{x})$. Second, the analysis state is also super-resolved using $\mathsf{F}$. The function $\mathsf{A}$ in Eq. (3a) is replaced with a composition $\mathsf{A}_{\text{HR}}=\mathsf{F}\circ\mathsf{A}_{\text{LR}}$ in Eq. (10a): the LR analysis is first calculated as $\bm{a}_{\text{LR}}=\mathsf{A}_{\text{LR}}(\bm{x},\bm{y})$ and then it is super-resolved as $\bm{a}_{\text{HR}}=\mathsf{F}(\bm{a}_{\text{LR}})$. The influences of these two changes are further discussed in Sections IV.2 and IV.3.
The ELBO is transformed by substituting Eqs. (10a)–(10c) into (2c), as in Section III.1:
$l_{\text{SRDA}}=\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y}))-\mathsf{F}(\bm{x})\right\rVert^{2}+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$ (13)
@@
$\bm{a}_{\text{HR}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y}))=\mathsf{F}(\mathsf{A}_{\text{LR}}(\bm{x},\bm{y})),$ (14a)
$\bm{a}_{\text{LR}}$ $=$ $\mathsf{A}_{\text{LR}}(\bm{x},\bm{y}),$ (14b)
$\bm{\hat{z}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})+\bm{\varepsilon}^{\text{T}}V(\bm{x},\bm{y})^{1/2}=\mathsf{F}(\bm{a}_{\text{LR}})+\sum_{i=1}^{n_{z}}\varepsilon_{i}\sqrt{v_{i}}.$ (14c)
@@
Eqs. (13) and (14c) corresponds to Eqs. (7) and (5) in Section III.1, respectively. Like in Eq. (7), the Monte Carlo method (i.e., the reparameterization trickKingma and Welling (2014)) is applied in Eq. (13) with the sample size being 1. In the right-hand side of Eq. (14c), the arguments of $\bm{x}$ and $\bm{y}$ are omitted, and the definitions of $\bm{a}_{\text{HR}}$ and $\bm{a}_{\text{LR}}$ are used. The first term of $l_{\text{SRDA}}$ represents the difference between the observation vector $\bm{y}$ and the estimated true state $\mathsf{H}(\bm{\hat{z}})$ in the observation space. The second term is the difference between the super-resolved analysis $\mathsf{F}(\bm{a}_{\text{LR}})$ and the super-resolved background $\mathsf{F}(\bm{x})$. The third term is a convex function of $v_{i}$, which takes the minimum when $v_{i}(\bm{x},\bm{y})=b$. These points indicate that the ELBO describes the balance between the observations and the super-resolved forecast in estimating the HR true state.
IV.2 Case of linear SR
@@
We interpret $l_{\text{SRDA}}$ in Eq. (13) using a linear SR operator and show that SR operators act as background error covariance matrices. Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning.Park, Park, and Kang (2003) When generation processes of LR images from HR images are linear (e.g., subsampling and blurring), the inverse processes, namely the SR processes, are also linear.Park, Park, and Kang (2003) Here, an SR operator is denoted as a matrix $F$, which has the size of $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) and maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors. Since SR is generally non-local, the matrix $F$ has non-zero off-diagonal elements.
@@
To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section III.2. The estimated true state $\bm{\hat{z}}$ in Eq. (14c) is then equal to $\bm{\hat{z}}=F\bm{a}_{\text{LR}}$. Further, the third term in Eq. (13) is ignored because it becomes constant.
Thus, $l_{\text{SRDA}}$ is reduced to
@@
$l_{\text{linear SRDA}}=\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(F\bm{a}_{\text{LR}})\right\rVert^{2}+\frac{1}{2b}\left(\bm{a}_{\text{LR}}-\bm{x}\right)^{\text{T}}\left(F^{\text{T}}F\right)\left(\bm{a}_{\text{LR}}-\bm{x}\right).$ (15)
@@
The matrix $F^{\text{T}}F$ is an $n_{x}\times n_{x}$ positive semi-definite matrix. Compared to the ELBO for 3D-Var in Eq. (8), $F^{\text{T}}F$ can be interpreted as the covariance matrix $B$ and describes the spatial correlations among the LR components.
When the observation operator is linear ($\mathsf{H}=H$) as in Section III.2, the minimizer for $l_{\text{linear SRDA}}$ is always obtained due to the convexity of $l_{\text{linear SRDA}}$. We first consider the case where $F^{\text{T}}F$ is positive definite (i.e., invertible). The minimizer is then derived, as in Eq. (9):
$\bm{a}_{\text{LR}}$ $=$ $\bm{x}+\widetilde{K}\left(\bm{y}-\widetilde{H}\bm{x}\right),$ (16a)
$\widetilde{K}$ $=$ $\widetilde{B}\widetilde{H}^{\text{T}}\left(rI_{n_{y}}+\widetilde{H}\widetilde{B}\widetilde{H}^{\text{T}}\right)^{-1},$ (16b)
where $\widetilde{B}=b\left(F^{\text{T}}F\right)^{-1}$ and $\widetilde{H}=HF$. The observation operator for the LR background is given by $\widetilde{H}=HF$ through the SR operator $F$. Generally, the Kalman gain matrix $\widetilde{K}$ has non-zero off-diagonal elements due to $F$ being non-diagonal. Thus, the LR background $\bm{x}$ is non-locally corrected by the observations $\bm{y}$, giving the LR analysis $\bm{a}_{\text{LR}}$ in Eq. (16a).
Even when $F^{\text{T}}F$ is positive semi-definite (i.e., at least one eigenvalue is zero), minimizers for $l_{\text{linear SRDA}}$ exist due to its convexity. In this case, a solution similar to Eq. (16a) is derived using pseudo-inverse matrices (not shown in detail).
IV.3 Case of non-linear SR
In practice, SR models are neural networks, so SR operators are non-linear. Since the nonlinear case includes the linear case, the non-local inference is achieved in both cases. Here, we briefly examine non-linear SR operators from a different perspective: the transformation of probability distributions.
We consider a complete non-linear SR operator $\mathsf{F}$, that is, a bijective SR operator. Rigorously speaking, SR is an inverse problem and SR operations are not bijective; however, a virtual bijective map is sometimes considered in theoretical studies of SR. In this case, any HR true state $\bm{z}$ has the corresponding LR state: $\bm{z}_{\text{LR}}=\mathsf{F}^{-1}(\bm{z})$ and $\bm{z}=\mathsf{F}(\bm{z}_{\text{LR}})$. We can regard the posterior and prior distributions in Eqs. (10a) and (10b) as non-Gaussian distributions for $\bm{z}_{\text{LR}}$:

### v4

@@ 3.2 Reduction of the ELBO to the 3D-Var objective function
where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from $\bm{x}$ and $\bm{y}$. This $\bm{a}$ is obtained by minimizing $l_{\text{3D-Var}}$. The first term of $l_{\text{3D-Var}}$, which comes from the reconstruction error in Eq. (2c), represents the difference between the observation vector $\bm{y}$ and the analysis $\bm{a}$ via the observation operator $\mathsf{H}$. The second term of $l_{\text{3D-Var}}$, which comes from the KL divergence in Eq. (2c), represents the difference between the analysis $\bm{a}$ and the background state $\bm{x}$. By minimizing $l_{\text{3D-Var}}$, $\bm{x}$ and $\bm{y}$ are integrated to estimate the true state as $\bm{a}$. Since $l_{\text{3D-Var}}$ is derived from the ELBO [i.e., $l_{\text{Gauss}}$ in Eq. (7)], the ELBO describes the balance between $\bm{x}$ and $\bm{y}$ in estimating the true state.
The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is a matrix $H$ [1].
@@
where $K$ is known as the Kalman gain matrix. These equations describe the non-locality of DA: an observation at one grid point is incorporated into the surrounding inferred values. This non-locality typically originates from the non-zero off-diagonal elements of the covariance matrix $B$. If $B$ is diagonal (i.e., all off-diagonal elements are zero), DA becomes local inference, and spatially sparse observations are incorporated into the inference only at their located grid points.
3.3 Implementation of DA using CVAE
DA can be performed by minimizing $l_{\text{Gauss}}$ in Eq. (7). However, this minimization at each assimilation time step would be computationally expensive. We propose an efficient approach to this minimization using a trained CVAE.
The loss function for VAEs, including CVAEs, is generally an ELBO [7, 8, 9, 11, 12]. This can be understood from the ELBO formula [Eq. (7)]. To evaluate $l_{\text{Gauss}}$, the analysis $\bm{a}$ and its covariance $V$ are first computed from given $\bm{x}$ and $\bm{y}$; the observation vector $\bm{\hat{y}}$ [$=\mathsf{H}(\bm{\hat{z}})$] is then reconstructed from $\bm{a}$ and $V$ via Eq. (5). The value of $l_{\text{Gauss}}$ decreases as $\bm{\hat{y}}$ becomes closer to the input $\bm{y}$. The architecture that reconstructs the input is a characteristic of autoencoders [10]. The second and third terms in Eq. (7) are the conditions for the estimated true state $\bm{\hat{z}}$. To incorporate these conditions, we use a CVAE, which is trained by minimizing the loss function $l_{\text{Gauss}}$. This training is unsupervised because $l_{\text{Gauss}}$ does not require the ground truth for $\bm{z}$. This unsupervised learning is achieved through the use of the ELBO because $l_{\text{Gauss}}$ is derived from the ELBO in Eq. (2c).
The CVAE consists of an encoder and a decoder (Fig. 2). The encoder estimates $\bm{a}$ and $V$ from the input $\bm{x}$ and $\bm{y}$, while the decoder reconstructs $\bm{\hat{y}}$. In other words, the encoder performs assimilation, while the decoder acts as the observation operator. The encoder learns how to integrate $\bm{x}$ and $\bm{y}$, yielding the analysis $\bm{a}$, while the decoder learns how to convert $\bm{a}$ into the observation space. Once the CVAE is trained, the encoder performs DA without minimizing $l_{\text{Gauss}}$ as long as the statistical properties of $\bm{x}$ and $\bm{y}$ do not change significantly.
@@ 3.3 Implementation of DA using CVAE
The covariance matrices $R$ and $B$ in $l_{\text{Gauss}}$ are treated as prescribed parameters (i.e., hyperparameters). Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices (i.e., having zero off-diagonal elements) [7, 8, 9, 11, 12]. However, the off-diagonal elements of $B$ are essential for the non-local inference in DA (Section 3.2). We adopt super-resolution (SR) to realize non-local inference using CVAEs, leading to a new theory of SRDA.
4 SRDA using the ELBO
4.1 Integration of SR into DA
To develop a theory of SRDA, we first specify the resolutions of a background state $\bm{x}$, observations $\bm{y}$, and the true state $\bm{z}$: $\bm{x}$ is defined at LR grid points, while $\bm{y}$ and $\bm{z}$ are defined at HR grid points. The vector forms are obtained by arranging three-dimensional grid point values into columns. Observations are not available at all locations due to various factors, such as measurement errors. These missing grid points are excluded when arranging observations into the column vector $\bm{y}$.
@@ 4.1 Integration of SR into DA
$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}$ (10a)
$\times$ $\exp\Bigg{\{}-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\Bigg{\}},$
@@
where each variance $v_{i}(\bm{x},\bm{y})$ ($i=1,\ldots,n_{z}$) is positive and determined by $\bm{x}$ and $\bm{y}$.
There are two essential changes from Eqs. (3) to (10). First, an SR operator $\mathsf{F}$ is introduced in Eq. (10b), mapping an $n_{x}$-dimensional vector into an $n_{z}$-dimensional vector. Since $\bm{x}$ and $\bm{z}$ are defined at LR and HR grid points, respectively (i.e., $n_{x}<n_{z}$), this mapping performs SR. Comparing Eqs. (3b) and (10b), the mean of the prior distribution $p\left(\bm{z}\mid\bm{x}\right)$ is replaced with the super-resolved background state $\mathsf{F}(\bm{x})$. Second, the analysis state is also super-resolved using $\mathsf{F}$. The function $\mathsf{A}$ in Eq. (3a) is replaced with a composition $\mathsf{A}_{\text{HR}}=\mathsf{F}\circ\mathsf{A}_{\text{LR}}$ in Eq. (10a): the LR analysis is first calculated and then super-resolved.
The ELBO for SRDA is obtained by substituting Eqs. (10a)–(10c) into Eq. (2c), as in Section 3.1:
$l_{\text{SRDA}}$ $=$ $\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})-\mathsf{F}(\bm{x})\right\rVert^{2}$ (13)
$\quad+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$
@@
$\bm{\hat{z}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})+\bm{\varepsilon}^{\text{T}}V(\bm{x},\bm{y})^{1/2}=\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}V^{1/2},$ (14a)
$\bm{a}_{\text{HR}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})=\mathsf{F}(\bm{a}_{\text{LR}}),$ (14b)
$\bm{a}_{\text{LR}}$ $=$ $\mathsf{A}_{\text{LR}}(\bm{x},\bm{y}),$ (14c)
@@
Eqs. (13) and (14a) correspond to Eqs. (7) and (5), respectively, where $V$ in Eq. (14a) is defined by Eq. (12). As in Eq. (7), the Monte Carlo method (i.e., the reparameterization trick [8]) is applied in Eq. (13) with the sample size being 1. In the right-hand side of Eq. (14a), the arguments of $\bm{x}$ and $\bm{y}$ are omitted, and the definition of $\bm{a}_{\text{HR}}$ [Eq. (14b)] is used. This $\bm{a}_{\text{HR}}$ is obtained by super-resolving $\bm{a}_{\text{LR}}$. The first term of $l_{\text{SRDA}}$ represents the difference between the observation vector $\bm{y}$ and the estimated true state $\mathsf{H}(\bm{\hat{z}})$ in the observation space. The second term is the difference between the super-resolved analysis $\bm{a}_{\text{HR}}$ [$=\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})$] and the super-resolved background $\mathsf{F}(\bm{x})$. The third term is a convex function of $v_{i}$ that takes the minimum when $v_{i}(\bm{x},\bm{y})=b$. The ELBO (i.e., $l_{\text{SRDA}}$) describes the balance between the observations and the super-resolved background state in estimating the HR true state.
4.2 Case of linear SR
The effects of non-locality of SR are examined separately for linear and nonlinear SR. For linear SR, we show that SR operators act as background error covariance matrices.
@@ 4.2 Case of linear SR
Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning [47]. When generation processes of LR images from HR images are linear (e.g., subsampling or blurring), the inverse processes, namely the SR processes, are also linear [47]. An SR operator is denoted by a matrix $F$ of size $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) that maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors.
@@
To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section 3.2. The estimated true state $\bm{\hat{z}}$ in Eq. (14a) is then approximated as $\bm{\hat{z}}=F\bm{a}_{\text{LR}}$. Further, the third term in Eq. (13) is ignored because it becomes constant.
Consequently, $l_{\text{SRDA}}$ is reduced to
@@
$l_{\text{SRDA-L}}=\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(F\bm{a}_{\text{LR}})\right\rVert^{2}+\frac{1}{2b}\left(\bm{a}_{\text{LR}}-\bm{x}\right)^{\text{T}}\left(F^{\text{T}}F\right)\left(\bm{a}_{\text{LR}}-\bm{x}\right),$ (15)
@@
where SRDA-L stands for SRDA with linear SR. The matrix $F^{\text{T}}F$ is an $n_{x}\times n_{x}$ positive semi-definite matrix. Compared to the ELBO for 3D-Var in Eq. (8), $F^{\text{T}}F$ acts as the covariance matrix $B$. Due to the non-locality of SR, $F^{\text{T}}F$ is non-diagonal and describes the spatial correlations among the LR components (Fig. 3).
Figure 3: Schematic of spatial correlation through a linear SR operator $F$. This $F$ super-resolves an LR image $\bm{w}$ to the HR image $\bm{\hat{w}}$ ($=F\bm{w}$). Here, $\bm{w}$ corresponds to $\bm{a}_{\text{LR}}-\bm{x}$ in Eq. (15). The two blue pixels in $\bm{w}$ contribute to the red pixel in $\bm{\hat{w}}$. By taking an inner product $\bm{\hat{w}}^{\text{T}}\bm{\cdot}\bm{\hat{w}}$, the two blue pixels are correlated with each other.
4.3 Case of nonlinear SR
In practice, SR operators are implemented as neural networks and are thus nonlinear. Since nonlinear SR encompasses linear SR, non-local inference is also achieved with nonlinear SR. In this section, non-locality is examined from a perspective of backpropagation through an SR neural network.
We consider the nonlinear version of Eq. (15):